# What If Moderation Didn’t Mean Suppression? A Case for Personalized Content Transformation

### v1

@@ 1. Introduction
These two scenarios highlight a fundamental challenge in online content moderation, defined by two interconnected problems. First, what counts as harmful varies drastically by user, over time, and across situations. Centralized, one-size-fits-all policies, designed to filter broadly proscribed content like hate speech or graphic violence, struggle to accommodate the deeply personal and contextual nature of content sensitivity (Roberts, 2019; Gillespie, 2018). Second, existing moderation tools, either those exercised by the platforms or given to users, rely on blunt interventions that effectively boil down to suppression of whole posts or accounts—whether through removal, algorithmic downranking that buries them out of sight, muting all posts containing certain keywords, or unfollowing. These interventions force users into the impossible trade-off Aisha and Rex face. These blunt instruments cannot differentiate between a single triggering element and the valuable surrounding content. The result is a choice between enduring harm and complete disconnection.
@@
We evaluate DIY-MOD through two user studies. Our first study (Section˜5), an in-situ evaluation of naturalistic browsing on Reddit feeds, finds that personalized transformation increases users’ sense of agency and enables them to safely engage with content they would otherwise avoid. Our second study (Section˜6), a controlled preference elicitation, identifies a set of principles for selecting effective content transformation, such as the need to provide cognitive closure. Together, our findings demonstrate that this approach offers a promising and user-valued alternative to suppressive content moderation. In summary, this paper makes the following contributions:
@@ 2.1. Centralized Moderation
Platform governance scholarship characterizes platforms as private rule-setters that curate and enforce boundaries for what their users may see (Gillespie, 2018). In practice, we refer to centralized moderation as arrangements where end-users have little to no say in defining or enforcing the rules (Sankin, 2017; Jahanbakhsh et al., 2022). Instead, decisions are made by platform policy teams, automated systems, or designated intermediaries such as community or instance administrators (Lee et al., 2024; Díaz and Hecht-Felella, 2021). While these intermediaries represent a form of delegated authority, their power is ultimately constrained by boundaries set by the platform. Platforms typically seek to avoid liability for third-party content (Keller, 2023a; Gillespie, 2010), but they increasingly act as setters of norms (Gillespie, 2018) and enforcers of the rules they establish.333 We take as given that clearly illegal content (e.g., CSAM (of Justice, [n.d.]), direct and credible incitement to violence) must be removed under law and platform policy. Our focus is lawful content that can nevertheless be distressing or harmful depending on the recipient and context. To operate at scale, these arrangements posit a common boundary of “harmful” content that can be expressed as rules and categories and applied uniformly.
@@
At platform scale, such a shared boundary needs to fit many norms and situations (Gillespie, 2020; Gorwa et al., 2020; Keller, 2023b). Fixed categories make enforcement easier and consistent but risk under-suppressing content that some decry as harmful, and over-suppressing content that some argue does not have the potential to harm (Gillespie, 2018; Lee et al., 2024). The underlying assumption in this approach is that content should be treated the same for everyone. But in practice, there is frequent and contentious debate on whether the policies themselves are legitimate or appropriate (Douek, 2021; Caplan, 2018).
@@
A universal boundary enables platforms to hand classification to AI systems and human moderators. Yet cases that sit at the edge, or events not covered by existing rules, produce inconsistency and ad hoc decisions (Myers West, 2018; Jahanbakhsh et al., 2023; Jahanbakhsh, 2023). Once flagged, enforcement usually means visibility limits such as downranking, gating, or removal (Gorwa et al., 2020). Softer measures like content warnings (CW) (Foa et al., 2007) still depend on centralized labels, where either the platform or the poster assigns a predefined category. Several scholars also propose applying restorative and transformative justice frameworks to platform governance as alternatives to these punitive approaches  (Hasinoff and Schneider, 2022; Xiao et al., 2023; Schoenebeck and Blackwell, 2020; Xiao et al., 2022)
@@
A limited form of decentralized decision making exists where subcommunities or instances set their own rules (Fiesler et al., 2018; Kuo et al., 2023), or when experiments like digital juries invite peers to deliberate on specific moderation cases (Fan and Zhang, 2020). Yet these approaches still require defining universal categories of acceptable content applied to all members, regardless of context or individual differences (Jahanbakhsh et al., 2022).
@@ 2.2. Harm as Situated and Subjective
Much debate has centered on where the boundary of harmful content lies. The pursuit of a universal boundary, however, is inherently flawed: it presumes that a single set of rules can capture the diversity of how people experience harm. We illustrate this with content usually seen as benign that can be profoundly harmful for individuals with specific lived experiences, such as phobias or PTSD triggers.
Specific phobias, some with genetic roots  (Kendler et al., 2001; Villafuerte and Burmeister, 2003), cause intense fear responses to otherwise low-risk stimuli (American Psychiatric Association, 2013; National Institute of Mental Health, 2022). Even digital exposure (Appleby, 2021; Michałowski et al., 2017) can provoke immediate physiological reactions, including panic attacks (Öst, 1989). These sensitivities can have societal consequences: people with trypanophobia (needle phobia) have reported that frequent needle imagery in vaccine news contributed to their vaccine hesitancy during COVID-19 (Appleby, 2021).
Trauma and life events also shape sensitivity. Each year, 3–4% of adults meet criteria for PTSD (Kessler et al., 2017). Pregnancy loss affects 10–20% of known pregnancies; content about pregnancy or infants–benign for many–can be acutely painful for some (Bardos et al., 2015). Eating disorders affect roughly 9% of the population and heighten sensitivity to food, weight, and body imagery (Hudson et al., 2007). In all cases, both the threshold and the interpretation of distress vary across people and over time.
Online environments make these differences especially visible. The same post can be celebratory for one person and painful for another, depending on needs and timing. Prior work on pregnancy loss documents this divergence in social media contexts (Andalibi and Forte, 2018; Andalibi, 2021). In eating-disorder contexts, people often seek support while trying to avoid content that exacerbates symptoms (Pater et al., 2016; Chancellor et al., 2016). A single category boundary cannot account for the range of experiences among users. A complementary lens from law and psychology also emphasizes subjective harm—the internal distress tied to a person’s experience, identity (Isham et al., 2023), or perception (Calo, 2011).
@@ 2.3. User Agency and Personalized Control
Architectural approaches: Middleware proposals (Fukuyama et al., 2021) ask platforms to expose interfaces so third parties can curate on a user’s behalf; users then can choose the provider whose values align with theirs. Federated and protocol-centric designs (e.g., Mastodon(Raman et al., 2019), Bluesky(Balduf et al., 2024)) push choice to the instance or client layer, where admins or client developers set policies and content boundaries (Masnick, 2019). In both cases, the gatekeeping moves closer to the end-users, but categories and thresholds are still defined upstream; end-users pick from presets.
User-facing tools: Prior HCI work shows strong demand for control over one’s feed (Jhaver et al., 2023; Jhaver and Zhang, 2025; Jahanbakhsh et al., 2022). Grassroots systems often layer community-generated safety signals directly onto existing platform interfaces. A prominent example is Shinigami Eyes: community labeling that marks spaces as welcoming or hostile to certain marginalized groups (Shinigami Eyes, 2024; DeVito et al., 2018; Sætra and Ese, 2023). Trust-based frameworks enable users to delegate judgments of content accuracy to selected trusted peers or services (Jahanbakhsh and Karger, 2024; Jahanbakhsh et al., 2022, 2023). At scale, these trust-based frameworks create collaborative signal pipelines that leverage the collective judgment of the community. Systems such as Crossmod (Chandrasekharan et al., 2019), help early detection of problematic comments that would be removed by moderators. These tools increase transparency and participation, but their mechanism remains suppressive of the entire content.
@@ 2.4. Towards Personalized Content Transformation
Status-quo moderation approaches operate at the post level, adjusting its visibility through removal, downranking, or limited distribution. These interventions affect the entire post regardless of whether it contains both harmful and valuable elements. User-side controls similarly filter exposure to entire posts or accounts through blocking, muting, or “show me less of this” feature.
We take a different path with content transformation, drawing on exposure therapy’s principle of safe engagement with weakened stimuli such as photos (Rothbaum and Schwartz, 2002; Yoshikawa et al., 2024). Prior HCI work has similarly modified visual content—through obfuscation (Li and Caine, 2022; Monteiro et al., 2024; Hasan et al., 2019) and cartoonification (Hasan et al., 2017)—to protect privacy while preserving utility. We apply this principle to minimize harm while maintaining the post’s informational and social value.
Recent AI advances provide the foundation for personalized transformation. Multimodal models (LLMs (Devlin et al., 2019; Kumar et al., 2024) and VLMs (Zhu et al., 2023; Liu et al., 2024)) distinguish central from peripheral content elements, enabling targeted interventions. Generative models then execture transformations: diffusion models remove or replace visual elements (Rombach et al., 2022), style-transfer changes presentation (Gatys et al., 2016), and language models can rewrite text while preserving meaning. Together, these tools allow selective modification of distressing elements while retaining core information. Because harm is subjective, the same piece of content can be transformed in several plausible ways. The challenge is choosing which best serves a given user. We adapt the LLM-as-a-judge approach (Zheng et al., 2023; Liu et al., 2023) to evaluate transformations through a personalized lens. The model receives both the transformed content and detailed user context to predict which intervention would best balance reducing harm with preserving information for that individual.
@@ 3.1. Methods
Our screening criteria focused on participants who could recall recent, concrete instances where otherwise ordinary content felt distressing because of personal context (e.g., phobias, trauma, troubling life experiences, or abuse), regardless of how platforms currently categorize or moderate such content. We recruited twelve participants whose experiences challenge conventional assumptions about harmful content. Their sensitivities spanned specific phobias (e.g., Kosmemophobia, Arachnophobia), trauma-related triggers (PTSD), and life events (e.g., pregnancy loss)—content ranging from what platforms ignore entirely to what they actively moderate. Following established practice of designing for ‘extra-ordinary’ users to improve systems for all (Pullin and Newell, 2007), we began with these cases. But we note that these experiences are not fundamentally different from what others may encounter. Content sensitivities can arise for anyone at different points in life. Upon completion of the full interview, each participant received a \$25 USD gift card.
@@ 3.3.1. A Spectrum of Triggers: From Phobias to Values
Beyond the relatively “easy” cases where restrictions are broadly accepted (Schoenebeck et al., 2023; Klonick, 2017), much of the tension over what should or should not be allowed in online spaces stems from the legitimate subjectivity of harm.
@@ 4. The System: DIY-MOD
We designed DIY-MOD as an open source browser extension555 The source code can be found at https://github.com/XXXXXXXX/diy_mod [anonymized for review] that creates a personalized content moderation layer between users and platform content. Consider Rex from our opening vignette. After installing DIY-MOD, he opens the extension popup. Through the conversational interface, he types: “I have an eating disorder that I’m trying to manage. Food pictures trigger cravings and obsessive thoughts I can’t control.” He then configures the filter with a high sensitivity level to apply only to images, setting the duration to ‘Always’.
@@
This scenario illustrates DIY-MOD’s three core components. First, users create filters through natural language conversation or image upload using the extension popup. Second, platform adapter intercepts incoming content during natural browsing before it is handed off for rendering (Piccardi et al., 2024; Kolluri et al., 2025). Third, our backend analyzes content using large vision-language models (LVLMs) and applies appropriate transformations and returns instructions and data to the adapter to render locally. We describe them in the following sections.
@@ 4.1. Filter Creation and Configuration
Users describe their sensitivities in natural language to DIY-MOD using their own words or by uploading example images. To capture the user’s specific needs, the system engages in conversational grounding (Cho and Rader, 2020; Brennan, 2014). The underlying language model is prompted to seek clarification when a description is ambiguous or underspecified, helping the user iteratively refine the filter. For instance, after a user uploads an image of jewelry, the system might ask: “Does this sensitivity apply to all jewelry or specific types?” Similarly, an initial filter for political content might prompt the system to ask: “Does this apply to all political topics, or specific ones like discussions of civilian casualties?”
@@
The extension popup interface provides essential controls within limited screen space. Users manage their filters (e.g., custom time limits, additional metadata, or editing filter descriptions) by accessing the options page (for Developers Team, 2012) of the extension. When creating filters, DIY-MOD can also detect potential overlaps and offer to modify existing filters.
@@ 4.2. Content Processing Client
The client employs platform-specific adapters that intercept network responses containing new content elements (Piccardi et al., 2024). This design keeps the filtering logic platform-independent, which we validated by implementing an adapter for Reddit. When adapters detect new posts, they extract text and image URLs before content is handed off for rendering. Only public content leaves the client and no user identifiers or profile information is transmitted to the server. The server analyzes this content against the user’s filters and returns personalized transformation instructions. The adapter then applies transformations to content within the post’s original layout before content becomes visible to the user. This client-side application of targeted modifications ensures transformations are applied at the recipient level (DG3) and preserves valuable surrounding elements within the content (DG4).

### v2

@@ 1. Introduction
These two scenarios highlight a fundamental challenge in online content moderation, defined by two interconnected problems. First, what counts as harmful varies drastically by user, over time, and across situations. Centralized, one-size-fits-all policies, designed to filter broadly proscribed content like hate speech or graphic violence, struggle to accommodate the deeply personal and contextual nature of content sensitivity (roberts2019; gillespie2018). Second, existing moderation tools, either those exercised by the platforms or given to users, rely on blunt interventions that effectively boil down to suppression of whole posts or accounts—whether through removal, algorithmic downranking that buries them out of sight, muting all posts containing certain keywords, or unfollowing. These interventions force users into the impossible trade-off Aisha and Rex face. These blunt instruments cannot differentiate between a single triggering element and the valuable surrounding content. The result is a choice between enduring harm and complete disconnection.
@@
We evaluate DIY-MOD through two user studies. Our first study (Section˜5), an in-situ evaluation of naturalistic browsing on Reddit feeds, finds that personalized transformation increases users’ sense of agency and enables them to safely engage with content they would otherwise avoid. However, because real-world browsing offers sporadic and unpredictable exposure to triggers, we conduct a second, controlled study (Section˜6) to compare transformation techniques. This preference elicitation identifies a set of principles for effective content transformation, such as the need to provide cognitive closure. Together, our findings demonstrate that this approach offers a promising and user-valued alternative to suppressive content moderation. In summary, this paper makes the following contributions:
@@ 2.1. Centralized Moderation
Platform governance scholarship characterizes platforms as private rule-setters that curate and enforce boundaries for what their users may see (gillespie2018). In practice, we refer to centralized moderation as arrangements where end-users have little to no say in defining or enforcing the rules (sankin2017activists; trustnet_user). Instead, decisions are made by platform policy teams, automated systems, or designated intermediaries such as community or instance administrators (lee2024people; diaz2021double). While these intermediaries represent a form of delegated authority, their power is ultimately constrained by boundaries set by the platform. Platforms typically seek to avoid liability for third-party content (keller2023carriage; gillespie2010politics), but they increasingly act as setters of norms (gillespie2018) and enforcers of the rules they establish.333 We take as given that clearly illegal content (e.g., CSAM (CSAM), direct and credible incitement to violence) must be removed under law and platform policy. Our focus is lawful content that can nevertheless be distressing or harmful depending on the recipient and context. To operate at scale, these arrangements posit a common boundary of “harmful” content that can be expressed as rules and categories and applied uniformly.
@@
At platform scale, such a shared boundary needs to fit many norms and situations (gillespie2020content; gorwa2020algorithmic; keller2023). Fixed categories make enforcement easier and consistent but risk under-suppressing content that some decry as harmful, and over-suppressing content that some argue does not have the potential to harm (gillespie2018; lee2024people). The underlying assumption in this approach is that content should be treated the same for everyone. But in practice, there is frequent and contentious debate on whether the policies themselves are legitimate or appropriate (douek2021governing; caplan2018content).
@@
A universal boundary enables platforms to hand classification to AI systems and human moderators. Yet cases that sit at the edge, or events not covered by existing rules, produce inconsistency and ad hoc decisions (myerswest2018; trustnet_ai; jahanbakhsh2023thesis). Once flagged, enforcement usually means visibility limits such as downranking, gating, or removal (gorwa2020algorithmic). Softer measures like content warnings (CW) (foa2007trigger) still depend on centralized labels, where either the platform or the poster assigns a predefined category. Several scholars also propose applying restorative and transformative justice frameworks to platform governance as alternatives to these punitive approaches  (hasinoff2022scalability; xiao2023addressing; schoenebeck2020reimagining; xiao2022sensemaking)
@@
A limited form of decentralized decision making exists where subcommunities or instances set their own rules (fiesler2018reddit; kuo2023unsung), or when experiments like digital juries invite peers to deliberate on specific moderation cases (fan2020digital). Yet these approaches still require defining universal categories of acceptable content applied to all members, regardless of context or individual differences (trustnet_user).
@@ 2.2. Harm as Situated and Subjective
Much debate has centered on where the boundary of harmful content lies. The pursuit of a universal boundary, however, is inherently flawed: it presumes that a single set of rules can capture the diversity of how people experience harm. We build this case by not focusing on the familiar “gray areas” that already generate public controversy (douek2021governing; caplan2018content; diaz2021double), but by turning to content that is rarely debated at all—content that most people overlook but that can be profoundly harmful for individuals with specific lived experiences, such as phobias or PTSD triggers.
Specific phobias, some with genetic roots  (kendler2001genetic; villafuerte2003genetic), cause intense fear responses to otherwise low-risk stimuli (apa2013dsm5; nimh2022phobias). Even digital exposure (kff_needle_phobia_article; michalowski2017set) can provoke immediate physiological reactions, including panic attacks (ost1989phobias). These sensitivities can have societal consequences: people with trypanophobia (needle phobia) have reported that frequent needle imagery in vaccine news contributed to their vaccine hesitancy during COVID-19 (kff_needle_phobia_article).
Trauma and life events also shape sensitivity. Each year, 3–4% of adults meet criteria for PTSD (kessler2017ptsd). Pregnancy loss affects 10–20% of known pregnancies; content about pregnancy or infants–benign for many–can be acutely painful for some (bardos2015miscarriage). Eating disorders affect roughly 9% of the population and heighten sensitivity to food, weight, and body imagery (hudson2007eating). In all cases, both the threshold and the interpretation of distress vary across people and over time.
Online environments make these differences especially visible. The same post can be celebratory for one person and painful for another, depending on needs and timing. Prior work on pregnancy loss documents this divergence in social media contexts (andalibi2018announcing; andalibi2021social). In eating-disorder contexts, people often seek support while trying to avoid content that exacerbates symptoms (pater2016hunger; chancellor2016quantifying). A single category boundary cannot account for the range of experiences among users. A complementary lens from law and psychology also emphasizes subjective harm—the internal distress tied to a person’s experience, identity (SHEEQ), or perception (calo2011boundaries).
@@ 2.3. User Agency and Personalized Control
Architectural approaches: Middleware proposals (fukuyama2021) ask platforms to expose interfaces so third parties can curate on a user’s behalf; users then can choose the provider whose values align with theirs. Federated and protocol-centric designs (e.g., Mastodon(mastodon), Bluesky(bluesky)) push choice to the instance or client layer, where admins or client developers set policies and content boundaries (masnick2019). In both cases, the gatekeeping moves closer to the end-users, but categories and thresholds are still defined upstream; end-users pick from presets.
User-facing tools: Prior HCI work shows strong demand for control over one’s feed (jhaver2023personalizing; jhaver2023users; trustnet_user). Grassroots systems often layer community-generated safety signals directly onto existing platform interfaces. A prominent example is Shinigami Eyes: community labeling that marks spaces as welcoming or hostile to certain marginalized groups (shinigami-eyes-2024; devito2023shinigami; saetra2023shinigami). Trust-based frameworks enable users to delegate judgments of content accuracy to selected trusted peers or services (trustnet_browser; trustnet_user; trustnet_ai). At scale, these trust-based frameworks create collaborative signal pipelines that leverage the collective judgment of the community. Systems such as Crossmod (eshwar_crossmods), help early detection of problematic comments that would be removed by moderators. These tools increase transparency and participation, but their mechanism remains suppressive of the entire content.
@@ 2.4. Towards Personalized Content Transformation
Status-quo moderation approaches operate at the post level, adjusting its visibility through removal, downranking, or limited distribution. These interventions affect the entire post regardless of whether it contains both harmful and valuable elements. User-side controls similarly filter exposure to entire posts or accounts through blocking, muting, or “show me less of this” feature. However, these controls have social costs. Blocking can appear hostile or signal ‘defeat’, which harassers treat as a badge of honor (heung2025ignorance). By severing ties completely (jhaver2018online), it forces a choice between enduring harm and losing social context and threat-monitoring capabilities.
We take a different path with content transformation, drawing on exposure therapy’s principle of safe engagement with weakened stimuli such as photos (rothbaum2002exposure; yoshikawa2024digital). Clinical variants, such as exposure and response prevention (ERP) for obsessive–compulsive disorder (bjorgvinsson2007obsessive; abramowitz2019exposure), use this principle in structured therapeutic programs. Recent work observes that triggers and tolerances both vary across people and change over time, and accordingly recommends that systems supporting ERP be adaptive (wang2025mentally). Prior HCI work has similarly modified visual content to protect privacy while preserving utility via obfuscation (li2022_obfuscation; monteiro2024_obfuscate; hasan2019_obfuscate) and cartoonification (hasan2017cartoonin_obfuscate). We apply this principle in a non-clinical, everyday setting: minimizing harm while maintaining the post’s informational and social value.
Recent AI advances provide the foundation for personalized transformation. Multimodal models (LLMs (devlin2019bert; kumar2023llm) and VLMs (zhu2023minigpt; liu2024improved)) distinguish central from peripheral content elements, enabling targeted interventions. Generative models then execture transformations: diffusion models remove or replace visual elements (rombach2022high-resolution-diffusion), style-transfer changes presentation (gatys2016image-style-transfer), and language models can rewrite text while preserving meaning. Together, these tools allow selective modification of distressing elements while retaining core information. Because harm is subjective, the same piece of content can be transformed in several plausible ways. The challenge is choosing which best serves a given user. We adapt the LLM-as-a-judge approach (zheng2023judging; liu2023gpteval) to evaluate transformations through a personalized lens. The model receives both the transformed content and detailed user context to predict which intervention would best balance reducing harm with preserving information for that individual.
@@ 3.1. Methods
Our screening criteria focused on participants who could recall recent, concrete instances where otherwise ordinary content felt distressing because of personal context (e.g., phobias, trauma, troubling life experiences, or abuse), regardless of how platforms currently categorize or moderate such content. We recruited twelve participants whose experiences challenge conventional assumptions about harmful content. Their sensitivities spanned specific phobias (e.g., Kosmemophobia, Arachnophobia), trauma-related triggers (PTSD), and life events (e.g., pregnancy loss)—content ranging from what platforms ignore entirely to what they actively moderate. Following established practice of designing for ‘extra-ordinary’ users to improve systems for all (extra_ordinary_user), we began with these cases. But we note that these experiences are not fundamentally different from what others may encounter. Content sensitivities can arise for anyone at different points in life. Upon completion of the full interview, each participant received a \$25 USD gift card.
@@ 3.3.1. A Spectrum of Triggers: From Phobias to Values
Beyond the relatively “easy” cases where restrictions are broadly accepted (schoenebeck2023online; klonick2017new), much of the tension over what should or should not be allowed in online spaces stems from the legitimate subjectivity of harm.
@@ 4. The System: DIY-MOD
We designed DIY-MOD as an open source browser extension 555 The source code can be found at https://github.com/UMichHCI/diy_mod that creates a personalized content moderation layer between users and platform content. Consider Rex from our opening vignette. After installing DIY-MOD, he opens the extension popup. Through the conversational interface, he types: “I have an eating disorder that I’m trying to manage. Food pictures trigger cravings and obsessive thoughts I can’t control.” He then configures the filter with a high sensitivity level to apply only to images, setting the duration to ‘Always’.
@@
This scenario illustrates DIY-MOD’s three core components. First, users create filters through natural language conversation or image upload using the extension popup. Second, platform adapter intercepts incoming content during natural browsing before it is handed off for rendering (piccardi2024reranking; kolluri2025alexandria). Third, our backend analyzes content using large vision-language models (LVLMs) and applies appropriate transformations and returns instructions and data to the adapter to render locally. We describe them in the following sections.
@@ 4.1. Filter Creation and Configuration
Users describe their sensitivities in natural language to DIY-MOD using their own words or by uploading example images. To capture the user’s specific needs, the system engages in conversational grounding (cho2020role; brennan2014grounding). The underlying language model is prompted to seek clarification when a description is ambiguous or underspecified, helping the user iteratively refine the filter. For instance, after a user uploads an image of jewelry, the system might ask: “Does this sensitivity apply to all jewelry or specific types?” Similarly, an initial filter for political content might prompt the system to ask: “Does this apply to all political topics, or specific ones like discussions of civilian casualties?”
@@
The extension popup interface provides essential controls within limited screen space. Users manage their filters (e.g., custom time limits, additional metadata, or editing filter descriptions) by accessing the options page (options_page) of the extension. When creating filters, DIY-MOD can also detect potential overlaps and offer to modify existing filters.
@@ 4.2. Content Processing Client
The client employs platform-specific adapters that intercept network responses containing new content elements (piccardi2024reranking). This design keeps the filtering logic platform-independent, which we validated by implementing an adapter for Reddit. When adapters detect new posts, they extract text and image URLs before content is handed off for rendering. Only public content leaves the client and no user identifiers or profile information is transmitted to the server. The server analyzes this content against the user’s filters and returns personalized transformation instructions. The adapter then applies transformations to content within the post’s original layout before content becomes visible to the user. This client-side application of targeted modifications ensures transformations are applied at the recipient level (DG3) and preserves valuable surrounding elements within the content (DG4).